# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

### v2

@@ 3.3 Hyperparametes
3.4 Models and Hardware
We conducted the experiments using closed and open models, the latest available at the time of running the experiments. Closed models (as they are often best performing, but less reproducible): OpenAI gpt-4-0125-preview (achiam2023gpt; openai2023gpt4turbo) with unconfirmed 1.76 trillion parameters (hereafter ‘GPT4-Turbo’) and gpt-4.1-2025-04-14 (openai2025gpt41docs; OpenAI_GPT-4.1_2025) with 1.8 trillion parameter (hereafter ‘GPT4.1’). Open models (more controlled and reproducible, even if not fully open): llama3.2:3b-instruct-q4_K_M (Dubey2024Llama3; meta2024llama), a quantized version of Meta Llama3.2 with 3 billion parameter (hereafter ‘Llama’) and qwen2.5:7b-instruct-q4_K_M (Yang2024Qwen2.5) with 7.61 billion parameters (hereafter ‘Qwen’). Closed model inference utilized the OpenAI API. Open model inference utilized Ollama (Ollama2023), as it is resource-efficient.
@@ 4 Results
Overall, our results set new state of the art for all four tasks, especially with GPT4-Turbo using Step-by-Step Prompting (except for Pado, where GPT4-Turbo tied with the highest baseline). See Table 2 and details below.
@@ 4.1 Closed Models
Both closed models (GPT4-Turbo and GPT4.1), when using Step-by-Step Prompting, outperformed the baselines by a large margin, except for Pado (where GPT4-Turbo tied with the highest baseline, and GPT4.1 was weaker than all baselines). With respect to the three axes, the models’ performance differed as follows:
@@
GPT4-Turbo demonstrated superior performance across the board and among all models, including GPT4.1, see Table 3. Step-by-Step Prompting experiments (Exp.3.x and 4.x) yielded the best scores, where absolute improvement over the strongest baseline was .23 for Fer-Loc, .22 for Fer-Ins, .23 for McRae, and a tie with Pado. Simple Prompting also outdid the baselines (except Pado).
GPT4.1 Step-by-Step Prompting also outdid all baselines except Pado, but curiously lower than GPT4-Turbo, in spite of being newer. Simple Prompting gave mixed results w.r.t. the baselines.
@@
Using Generated Sentences did worse or similar to using Lemma Tuples in all tasks and LLMs, except Pado with GPT4-Turbo.
@@
Using predefined categories compared to numeric output (Exp.x.2 vs. Exp.x.1) had zero, or small mixed effect, contrary to our hypothesis.
With GPT4-Turbo, using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) leads to a mixed effect across the board. Regardless of input and reasoning forms, the categorical score is lower than the numerical score in Fer-Loc. For Fer-Ins and Pado, the categorical score is either the same or higher than the numerical score, as the gain in Fer-Ins was .04 in Exp2 and in Pado was .04 in Exp3. Using categorical output in McRae lowers the score in general, except for Exp4, where the gain was .02.
@@ 4.2 Open Models
The performance of the open models was lower than the closed models, and lower than several baselines across the board, except for Fer-Loc using Step-by-Step Prompting with Llama and McRae with Qwen (mostly in Exp.4.x. Details are below:
@@
Using Generated Sentences instead of Lemma Tuples had mostly no effect with Llama, except for gains in McRae. With Qwen there were mostly gains except for lower scores in Exp.4.x in FerLoc.
@@ 5 Discussion
The core questions we aim to address in this work are: Do LLMs have the kind of linguistic knowledge required for thematic fit estimation? If so, (Q1) does Step-by-Step Prompting help utilize the model’s internal linguistics knowledge for this task? (Q2) Would providing more context by including sentences as part of the input help? (as opposed to providing only lemma tuples with semantic roles) (Q3) Do predefined thematic fit output categories better elicit the model’s internal linguistic knowledge compared to numeric output?
Overall, we conclude that closed models acquired much of the linguistic knowledge required for thematic fit estimation. Results with both GPT models on each of the test sets achieved or even surpassed all previous baselines, which were trained on labeled linguistic data (even if not directly trained for thematic fit estimation). This is impressive,999The available inter-annotator agreement or correlation information for our testsets indicates the upper bound for these tasks is between the mid-60’s and mid-70’s. GPT4-Turbo’s scores got close, but no single setting was always the best. assuming closed models were not trained on this task (sadly, we cannot know for sure because they are, well, closed). Next, we discuss the findings by the three axes (as defined in §3) and then move to analyzing additional aspects.
@@
Why was there a difference in the Step-by-Step Prompting effect between the closed and open models? Step-by-Step Prompting is designed to help the models reason by first eliciting useful properties of arguments and roles, before the final thematic fit estimation. However, our analysis on a sample of random tuples revealed that open models are weak at inferring the essential properties that are most expected for an argument, to serve well in a specific semantic role. Input from this step, containing invalid or irrelevant properties, often derailed the final reasoning output quality (See Appendix §E).
Why did GPT4-Turbo outperform GPT4.1 in spite of being older? We observed that GPT4.1’s ratings were often more ‘cautious’, meaning it rarely output very high or very low scores. We verified this by calculating the variance of each model’s output for each task and experimental setting (Appendix §B, Tables 7-8), finding a negative correlation between the variance and the correlation score (). The lower variance of GPT4.1 means it is more likely to do worse on clear-cut inputs (with very high or very low thematic fit).
@@
In principle, providing more context enhances models’ ability to comprehend complex tasks. This was generally reflected in our experiments with open models, but not closed models (with few exceptions).
Why do generated sentences only help open models? Although the open models did not outperform the closed models, they often better leveraged the input with generated sentences, and improved the correlation score () relative to the tuple input. But sentence input is likely to outperform tuple input only if these sentences are “good” (semantically coherent, with the given argument in the given role for the given predicate). Moreover, “bad” sentences are likely to derail the models’ performance. Therefore, we analyzed a random sample of 20 tuples, 5 generated sentences for each, altogether 100 sentences per model. We found that the open Qwen did better in filtering out “bad” sentences (in spite of having fewer parameters): its semantic filtering let through no False Positives (FPs), thus successfully leveraging the sentence inputs. In contrast, the closed GPT4-Turbo let through 6 FPs (Table 6). We saw similar results comparing filtering by Qwen vs. GPT4-Turbo, both on GPT4-Turbo’s sentences.101010Note that Qwen had higher False Negative rate, but this would only result in regressing to tuple input (§3.1 Axis 2). With it, closed models clearly generated more “good” sentences (80 vs. 60 True Positives + False Negatives in our sample). Therefore we caution against jumping into conclusions which model gained more relevant knowledge.
@@
There are arguments in favor of either categorical and numeric output. On the one hand, we expect predefined output categories to yield better results than numeric output, because LLMs are primarily designed for text-based tasks. On the other hand, in spite of this, LLMs have also shown impressive performance in math (even if not perfect). Also, the ground-truth thematic fit scores were averaged across the participants, providing a graded scale, and therefore perhaps a numeric output would better correlate with that. Indeed, output form had mostly small-to-zero effect (up to 0.02 difference) for closed models, but categorical output gave mixed results, mostly worse than numeric output in open models. We hypothesize this may be due to the differences in post-training between the closed and open models, but we currently have no solid explanation for this difference, and we leave this for future work.
@@ 6 Conclusion and Future Work
We set out to discover if, or to what extent, autoregressive LLMs possess linguistic and other knowledge required to estimate the thematic fit of event participants. The tested LLMs set new state-of-the-art on each test set (and tied on Pado). Still, our takeaway is that thematic fit is not yet “solved”, in spite of this clear progress. Closed models111111We observed that GPT4.1 did worse than GPT4-Turbo, in spite of being newer, likely because 4.1 is too ‘cautious’ (§5). yielded higher results than open models, indicating that closed models, especially GPT4-Turbo, acquired much of the linguistic knowledge required for thematic fit estimation. However, no single axis setting consistently yielded the highest scores. Moreover, open models where better at filtering “bad” generated sentences, so we caution against rushing to interpret the results as showing that open models gained less relevant knowledge. It is still unclear if models need additional knowledge, or if the right elicitation has not been found yet.
@@
(2) Generated sentences hurt closed models; We saw LLMs struggle with generating sentences where the given argument is a better fit to a semantic role different than the given one. To tackle this, we used novel semantic filtering, which, to our surprise, worked better for open models, because the latter were better at filtering out False Positives.
(3) Predefined categorical output had a mixed effect (relative to numeric output): zero-to-small effect with closed models and more negative effect with open models.
In future work, we hope to improve Step-by-Step Prompting, following our analysis, especially by generating better sentences as context, and better filtering out “bad” sentences. If successful, LLMs may also be used to study issues in thematic fit, in line with futrell2025how, BabyLM efforts121212https://babylm.github.io, and others who believe that LLMs can be useful tools to study linguistic phenomena.
@@
LLMs can be very sensitive to prompt phrasing, so it is possible that different prompts may yield very different results. Also, prompt engineering is tricky; thus, tweaking the prompts may lead to different results. The richness of prompt variation (prompt engineering) depends on human resources (different people use language differently) and funding/time resources, all of which were limited. Future work may achieve better results with more resources, if available. Due to the scarcity of evaluation data, we could not use prompt optimization methods such as DSPy Khattab2023DSPyCD or fine-tune the models.
@@ Appendix A Prompt Design
Each experiment has a specific input, output, and reasoning design, as detailed in Sec. §3.1. The corresponding prompt design for each experiment is specified here.
A.1 Simple Prompting with Head Lemma Tuples (Exp.1.1 — Exp.1.2)
A simple thematic fit prompt is used in the Simple Prompting with Head Lemma Tuples setting. It is constructed from a single prompt template that directly asks for the thematic fit score of a predicate, argument, role tuple. Here is an example of the filled template:
@@
As mentioned above, there are two output settings: numeric score (Exp.1.1) or categorical score (Exp.1.2). We extend the simple prompt as follows:
@@
A.2 Semantic Filtering Process for Generated Sentences
@@
A.3 Simple Prompting with Generated Sentences (Exp.2.1 — Exp.2.2)
The difference between experiments 1.x and 2.x is the input type. We augment the simple prompt with a natural language sentence previously generated by the model in the ‘sentence generation’ phase, from the same lemma tuples as in Exp.1.x:
@@
A.4 Step-by-Step Prompting with Lemma Tuples (Exp.3.1 — Exp.3.2)
Experiments 3.x are the enhanced version of the baseline Exp.1.x, where the reasoning steps precede the final semantic-fitness prompt. The reasoning phase is done through Step-by-Step Prompting, where a series of four prompts helps reveal the reasons/internal knowledge that the model relies on to rate the thematic fit of the given data. These four prompts are:
What salient essential properties or core characteristics best describe the essence of ‘pizza’?131313In an early version of this experiment, only the semantic role name was mentioned. After analyzing the early version, we suspected that the LLMs may have not made the connection between roles such as ‘Arg1’ and their PropBank sense in this context. Therefore. We added the word ‘PropBank’ before the semantic role in the Pado and McRae datasets, as their semantic roles were PropBank-based, unlike the Ferretti datasets. This indeed improved the results on these sets by 2-3% (absolute).
@@
A.5 Step-by-Step Prompting with Generated Sentences (Exp.4.1—Exp.4.2)
This last set of experiments combines Step-by-Step Prompting in Exp.3.x and the generated sentence input type as in Exp.2.x. In the later version of this experiment, the first four prompts were as in Exp.3.x but with the addition of the generated sentence in each prompt. Then the simple thematic fit prompt follows, also with one of the generated sentences as context, as in Exp.2.x :
@@
Appendix B Experimental Results with Variance
Table 7 and 8 show the Spearman correlation scores along with the variance for each experiment across the board. These tables support our explanation for the difference in performance between GPT4-Turbo and GPT4.1.
| Fer-Loc | Fer-Ins | Pado | McRae | |||||
| BG: GSD2015 | .29 | - | .42 | - | .53 | - | .36 | - |
| B0: NN RF | .44 | - | .45 | - | .52 | - | .38 | - |
| B1: ResRoFA-MT | .46 | - | .48 | - | .53 | - | .43 | - |
| B2: 20% subset v2 | - | - | - | - | .43 | - | .44 | - |
| B3a: Glove-shared-not-tuned | .30 | - | .34 | - | .49 | - | .31 | - |
| B3b: Glove-shared-tuned | .29 | - | .29 | - | .53 | - | .33 | - |
| B3c: Glove-shared-tuned with 20%M | .35 | - | .43 | - | .59 | - | .43 | - |
| Exp.1.1 | .63 | .121 | .64 | .132 | .35 | .168 | .54 | .159 |
| Exp.1.2 | .61 | .081 | .65 | .120 | .40 | .142 | .58 | .131 |
| Exp.2.1 | .48 | .085 | .49 | .075 | .42 | .149 | .55 | .136 |
| Exp.2.2 | .46 | .063 | .54 | .081 | .45 | .154 | .56 | .130 |
| Exp.3.1 | .69 | .097 | .70 | .115 | .48 | .125 | .65 | .137 |
| Exp.3.2 | .68 | .108 | .68 | .132 | .46 | .118 | .66 | .138 |
| Exp.4.1 | .57 | .077 | .66 | .074 | .58 | .130 | .58 | .133 |
| Exp.4.2 | .58 | .091 | .66 | .092 | .59 | .132 | .59 | .145 |
| Fer-Loc | Fer-Ins | Pado | McRae | |||||
| BG: GSD2015 | .29 | - | .42 | - | .53 | - | .36 | - |
| B0: NN RF | .44 | - | .45 | - | .52 | - | .38 | - |
| B1: ResRoFA-MT | .46 | - | .48 | - | .53 | - | .43 | - |
| B2: 20% subset v2 | - | - | - | - | .43 | - | .44 | - |
| B3a: Glove-shared-not-tuned | .30 | - | .34 | - | .49 | - | .31 | - |
| B3b: Glove-shared-tuned | .29 | - | .29 | - | .53 | - | .33 | - |
| B3c: Glove-shared-tuned with 20%M | .35 | - | .43 | - | .59 | - | .43 | - |
| Exp.1.1 | .51 | .046 | .56 | .105 | .40 | .145 | .64 | .158 |
| Exp.1.2 | .50 | .051 | .56 | .105 | .40 | .147 | .58 | .149 |
| Exp.2.1 | .32 | .024 | .33 | .050 | .33 | .096 | .40 | .058 |
| Exp.2.2 | .26 | .027 | .37 | .061 | .33 | .114 | .37 | .066 |
| Exp.3.1 | .59 | .042 | .59 | .082 | .45 | .111 | .64 | .118 |
| Exp.3.2 | .56 | .062 | .59 | .097 | .42 | .127 | .63 | .136 |
| Exp.4.1 | .56 | .030 | .53 | .056 | .42 | .104 | .44 | .061 |
| Exp.4.2 | .53 | .043 | .53 | .071 | .42 | .121 | .46 | .074 |
Appendix C Earlier round of Experiment

### v3

@@ 3.3 Hyperparameters
3.4 LLMs
We conducted the experiments using closed and open models, the latest available at the time of running the experiments. Closed models (as they are often best performing, but less reproducible): OpenAI gpt-4-0125-preview (Achiam et al., 2023; OpenAI, 2023) with unconfirmed 1.76 trillion parameters (hereafter ‘GPT4-Turbo’) and gpt-4.1-2025-04-14 (OpenAI, 2025a, b) with 1.8 trillion parameter (hereafter ‘GPT4.1’). Open models (more controlled and reproducible, even if not fully open): llama3.2:3b-instruct-q4_K_M (Dubey et al., 2024; Meta, 2024), a quantized version of Meta Llama3.2 with 3 billion parameter (hereafter ‘Llama’) and qwen2.5:7b-instruct-q4_K_M (Yang et al., 2024) with 7.61 billion parameters (hereafter ‘Qwen’). Closed model inference utilized the OpenAI API. Open model inference utilized Ollama (Ollama Developers, 2023).
@@ 4.1 Closed Models
Overall, our results (Table 2) set a new state-of-the-art for all four tasks, especially with GPT4-Turbo using Step-by-Step PromptingḂoth GPT4-Turbo and GPT4.1 outperformed the baselines by a large margin, except for Pado (where GPT4-Turbo tied with the highest baseline, and GPT4.1 was weaker than most baselines). GPT4-Turbo demonstrated superior performance across the board, outperforming also the newer GPT4.1. With respect to the three axes, the model performance differed as follows:
@@
Simple Prompting with GPT4-Turbo outdid all baselines (except Pado), while results with GPT4.1 were mixed, in spite of this model being newer.
For both GPT4-Turbo and GPT4.1, Step-by-Step Prompting (Exp.3.x and 4.x) yielded the best scores, demonstrating the effectiveness of this approach. For GPT4-Turbo, absolute improvement over the strongest baseline was for Fer-Loc, for Fer-Ins, for McRae, and a tie with Pado.
@@
For both closed models, using Generated Sentences (Exp.2.x and 4.x) rather than Lemma Tuples was detrimental. The only exception was for Pado with GPT4-Turbo, where using this approach together with Step-by-Step Prompting resulted in our highest score for this test set.
@@
Using predefined categories compared to numeric output (Exp.x.2 vs. Exp.x.1) had zero, or small mixed effect, contrary to our hypothesis. With GPT4-Turbo, using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) leads to a mixed effect across the board. Regardless of input and reasoning forms, the categorical score is lower than the numerical score in Fer-Loc. For Fer-Ins and Pado, the categorical score is either the same or higher than the numerical score, as the gain in Fer-Ins was .04 in Exp2 and in Pado was .04 in Exp3. Using categorical output in McRae lowers the score in general, except for Exp4, where the gain was .02.
@@ 4.2 Open Models
The performance of the open models was lower than the closed models, and lower than several baselines across the board, except for Fer-Loc using Step-by-Step Prompting with Llama and McRae with Qwen (mostly in Exp.4.x). Details are below:
@@
Using Generated Sentences (Exp.2.x and 4.x) instead of Lemma Tuples had mostly no effect with Llama, except for gains in McRae. With Qwen there were mostly gains except for lower scores in Exp.4.x in FerLoc.
@@ 5 Discussion
The core questions we aim to address in this work are: Do LLMs have the kind of linguistic knowledge required for TFE? If so, (Q1) does Step-by-Step Prompting help utilize the model’s internal linguistics knowledge for this task? (Q2) Would providing more context by adding sentences to the input help? (as opposed to providing only lemma tuples and roles) (Q3) Do predefined thematic fit output categories better elicit the model’s internal linguistic knowledge compared to numeric output?
Overall, we conclude that closed models acquired much of the linguistic knowledge required for TFE. Results with both GPT models on each of the test sets achieved or even surpassed all previous baselines, which were trained on labeled linguistic data (even if not directly annotated for TFE, such as SRL). This is impressive,888The available inter-annotator agreement or correlation information for our testsets indicates the upper bound for these tasks is between the mid-60’s and mid-70’s. GPT4-Turbo’s scores got close, but no single setting was always the best. assuming closed models were not trained on this task (sadly, we cannot know for sure because they are, well, closed). Next, we discuss the findings by the three axes (as defined in §3) and then move to analyzing additional aspects.
@@
Why was there a difference in the Step-by-Step Prompting effect between the closed and open models? Step-by-Step Prompting is designed to help the models reason by first eliciting useful properties of arguments and roles, before the final TFE. However, our analysis on a sample of random tuples revealed that open models are weak at inferring the essential properties that are most expected for an argument, to serve well in a specific semantic role. Input from this step, containing invalid or irrelevant properties, often derailed the final reasoning output quality. Moreover, if a reasoning step failed early on, it often derailed the rest of the chain (see Figure 2) which may explain the score losses in some of the experiments.
Why did GPT4-Turbo outperform GPT4.1 in spite of being older? We observed that GPT4.1’s ratings were often more ‘cautious’, meaning it rarely output very high or very low scores. We verified this by calculating the variance of each model’s output for each task and experimental setting, finding a negative correlation between the variance and the correlation score (). The lower variance of GPT4.1 means it is more likely to do worse on inputs with very high or very low thematic fit.
@@
In principle, providing more context enhances the models’ ability to comprehend complex tasks. This was generally reflected in our experiments with open models, but not closed models (with few exceptions).
Why do generated sentences only help open models? Although the open models did not outperform the closed models, they often better leveraged the input with generated sentences, and improved the correlation score () relative to the tuple input. But sentence input is likely to outperform tuple input only if these sentences are “good” (semantically coherent, with the given argument in the given role for the given predicate). Moreover, “bad” sentences are likely to derail the models’ performance. Therefore, we analyzed a random sample of 20 tuples, 5 generated sentences for each, altogether 100 sentences per model. We found that the open Qwen did better in filtering out “bad” sentences (in spite of having fewer parameters): its semantic filtering let through no False Positives (FPs), thus successfully leveraging the sentence inputs. In contrast, the closed GPT4-Turbo let through 6 FPs (Table 6). We saw similar results comparing filtering by Qwen vs. GPT4-Turbo, both on GPT4-Turbo’s sentences.999Note that Qwen had higher False Negative rate, but this would only result in regressing to tuple input (§3.1 Axis 2). With it, closed models clearly generated more “good” sentences (80 vs. 60 True Positives + False Negatives in our sample). Therefore, we caution against jumping to conclusions which model gained more relevant knowledge.
@@
There are arguments in favor of either categorical and numeric output. LLMs are primarily designed for text-based tasks, so we expect predefined output categories to yield better results than numeric output. However, LLMs have also shown impressive performance in math (even if not perfect). Also, the ground-truth thematic fit scores were averaged across the participants, providing a graded scale, and therefore perhaps a numeric output would better correlate with that. Indeed, output form had almost no effect for closed models, but categorical output gave mostly worse results with open models. We hypothesize this may be due to the differences in post-training between the closed and open models, but we currently have no solid explanation for this difference, and we leave this for future work.
@@ 6 Conclusion and Future Work
We set out to discover if, or to what extent, autoregressive LLMs possess linguistic and other knowledge required to estimate the thematic fit of event participants. The tested LLMs set new state-of-the-art on each test set (and tied on Pado). In spite of this clear progress, we conclude that TFE is not yet “solved”. Closed models101010We observed that GPT4.1 did worse than GPT4-Turbo, in spite of being newer, likely because 4.1 is too ‘cautious’ (§5). yielded higher results than open models, indicating that closed models, especially GPT4-Turbo, acquired much of the linguistic knowledge required for TFE. However, no single axis setting consistently yielded the highest scores. Moreover, open models were better at filtering “bad” generated sentences, so we caution against rushing to interpret the results as showing that open models gained less relevant knowledge. It is still unclear if models need additional knowledge or a different elicitation.
@@
(2) Generated sentences hurt closed models; We saw LLMs struggle with generating sentences where the given argument is a better fit to a semantic role different than the given one. To tackle this, we used novel semantic filtering, which, to our surprise, worked better for open models, because they were better at filtering out False Positives.
(3) Predefined categorical output (vs. numeric output) had zero-to-small effect with closed models and more negative effect with open models.
In future work, we hope to improve Step-by-Step Prompting, following our analysis, especially by generating better sentences as context, and better filtering of “bad” sentences. If successful, LLMs may also be used to study linguistic issues in thematic fit, in line with Futrell and Mahowald (2025), BabyLM efforts111111https://babylm.github.io, and others who believe that LLMs can be useful tools to study linguistic phenomena.
@@
LLMs can be very sensitive to prompt phrasing, so it is possible that different prompts may yield very different results. Also, prompt engineering is tricky; thus, tweaking the prompts may lead to different results. The richness of prompt variation (prompt engineering) depends on human resources (different people use language differently) and funding/time resources, all of which were limited. Future work may achieve better results with more resources, if available. Due to the scarcity of evaluation data, we could not use prompt optimization methods such as DSPy Khattab et al. (2024) or fine-tune the models.
@@ Appendix A Prompt Design
Each experiment specified in §3.1 has a prompt design with specific input, output, and reasoning:
Simple Prompting with Lemma Tuples (Exp.1.1–Exp.1.2)
A single prompt template (‘simple thematic fit’) that directly asks for the thematic fit score of a predicate, argument, role tuple (Fig. 1).
@@
We extend the prompt with two output settings:
@@
Sentence Generation and Semantic Filtering
@@
Simple Prompting with Generated Sentences (Exp.2.1–Exp.2.2)
We augment the prompt in Exp.1.x with sentences generated by the model from the same lemma tuples:
@@
Step-by-Step Prompting with Lemma Tuples (Exp.3.1–Exp.3.2)
Experiments 3.x are the enhanced version of Exp.1.x, where we break down the thematic fit task into a series of 4 prompts (‘Step-by-Step Prompting’). See Fig. 1.
What salient essential properties or core characteristics best describe the essence of ‘pizza’?121212In an early version of this experiment, only the semantic role name was mentioned. After analyzing it, we suspected that the LLMs may have not made the connection between roles such as ‘Arg1’ and their PropBank sense in this context. Therefore we added the word ‘PropBank’ before the semantic role for Pado and McRae, as their semantic roles were PropBank-based, unlike the Ferretti datasets. This indeed improved the results on these sets by 2-3% (absolute).
@@
Step-by-Step Prompting with Generated Sentences (Exp.4.1–Exp.4.2)
This experiment uses Step-by-Step Prompting as in Exp.3.x but with the addition of the generated sentence as in Exp.2.x:
@@
Appendix B Earlier round of Experiment